3x+4y=-8 and -5x+y=-25
Found 2 solutions by stanbon, RAY100:
Answer by stanbon(75887)

3x+4y=-8
-5x+y=-25
----
Multiply thru the 2nd equation by 4 to prepare for elimination:
----
3x+4y = -8
-20x+4y = -100
-------------------------
Subtract the 2nd from the 1st and solve for "x":
23x = 92
x = 4
---
Substitute into 3x+4y=-8 to solve for "y":
3*4 + 4y = -8
3 + y = -2
y = -5
------------------
Cheers,
Stan H.

Answer by RAY100(1637)

(1) 3x+4y = -8
.
(2) -5x +y = -25,,,,,or y= 25+5x,,,,,
.
subst in (1),,,,3x +4(25+5x) = -8
.
3x +100+20x = -8
.
23x = -108
.
x= -4.7,,,,,,,y=25+5(-4.7) = 1.5
.
.
check,,,(1) 3(-4.7) +4(1.5) = -8
.
-14.1 +6 = -8,,,,,ok
.
(2) -5(-4.7) + 1.5 = -25,,,ok
.